# Variability in Footpoint Mapping of Bursty Bulk Flows Using Tsyganenko Models:

### v2

@@ Abstract
Magnetospheric-ionospheric coupling studies often rely on multi-spacecraft conjunctions, which require accurate magnetic field mapping tools. For example, linking measurements from the magnetotail with those in the ionosphere involves determining when the orbital magnetic footpoint of THEMIS or MMS intersects with the footpoint of Swarm. One of the most commonly used tools for tracing magnetic field lines are the Tsyganenko models. In this study, we aim to analyze how the footpoint locations are impacted by the inputs parameters of these models, including solar wind conditions, geomagnetic activity, and the location in the magnetotail. A dataset of 2394 bursty bulk flows (BBFs) detected by MMS was mapped to Earth’s ionosphere with four different Tsyganenko models. The MLT position showed a variability of approximately ±$\pm 1$ hour MLT across the models. Footpoint locations were linked to the dawn-dusk position of the BBFs, with differences between models associated with variations in the interplanetary magnetic field clock angle. The MLAT values exhibited greater variability of approximately ±$\pm 2^{\circ}$ around the mean value, with a systematic shift toward higher latitudes observed in two of the four models. This position is also influenced by the input parameters of the model representing the dynamics of Earth’s magnetosphere, where stronger magnetospheric activity typically corresponds to lower latitudes. The uncertainty on the BBF footpoint location impacts the number of conjunctions with Swarm. Generally, Swarm B exhibited more conjunctions than Swarm A or C in the northern hemisphere. However, when considering only Swarm-BBF conjunctions where the distance between footpoints computed with T89 and T96 is smaller than the BBF footprint size, the number of conjunctions is reduced to less than half of the total.
@@
Bursty bulk flows footpoint – Tsyganenko mapping – Field-aligned currents
@@ 1 Introduction
Although model accuracy (i.e., how closely the derived footpoint matches its true value) is not the focus of this study, previous research has highlighted limitations in these models. For instance, Pulkkinen and Tsyganenko (1996) analyzed the accuracy of the T89 model, finding that mapping errors were largest on the dayside, while on the nightside they ranged from $0.5\degree$ to $1.8\degree$ in latitude (63°72°63°72°63$\degree$–72$\degree$63 ° – 72 °), increasing with geomagnetic activity. Similarly, Shevchenko et al. (2010) reported errors of $\sim 1\degree$ when evaluating the accuracy of the T96 model. Thus, the T96 model performs a better mapping than T89, but in both cases, the mapping error increases with magnetic activity. These discrepancies have been observed in case studies. For instance, Opgenoorth et al. (1994) found that a localized FAC was observed a few degrees north compared to the location obtained with the T89 model. Similarly, Wei et al. (2021) noted differences in the ground track mapping when using three different Tsyganenko models (T87, T89, and T96) while investigating the characteristics and responses of the magnetosphere-ionosphere-ground system during a case study.
This study utilizes a large database of BBF events observed by the Magnetospheric Multiscale (MMS) spacecraft (Burch et al., 2016) to statistically compare the mapped footpoint locations among four different Tsyganenko models. We refer to the footpoint as a unique point representing the BBF and do not consider the spatial size of the BBF mapped into the ionosphere (footprint). In this work, we do not address the error in footpoint location, since this implies establishing a ground truth for the correct footpoint for each event, which is beyond the scope of this analysis. Instead, we analyze how the footpoint locations are impacted by the inputs parameters of these models, including solar wind conditions, geomagnetic activity, the location of the BBF in the magnetotail, and the tilt angle. We report how the number of conjunctions between BBFs and Swarm spacecraft changes when the conjunction definition is adjusted. Specifically, we examine how increasing the time interval between the detection of a BBF and the Swarm satellites’ path through/nearby the BBF footpoint, as well as varying the footpoint magnetic latitude error, affects the occurrence and quality of conjunctions. Although we did not conduct a study on the error in the location, we hope that the quality indicator will help researches to determine when to be more cautious when studying conjunctions.
@@ 2.1 The BBF database
Since our goal is to find the ground footpoint of the BBFs, we have restricted our database to Earthward-moving BBFs, which constitute 89% of the total database. The mean and median duration of these Earthward BBFs are 3.60 minutes and 2.63 minutes, respectively, with approximately 80% of the events lasting less than 5 minutes. After excluding an outlier with a duration of $\approx 3~{}$hours, the BBF dataset consists of 2135 events, with the longest event lasting 35 minutes.
@@
Additionally, Figure 1b-c shows the MMS orbit for the second MMS season (2018-05-25 to 2018-09-28) and the BBFs detected during the same period (blue dots correspond BBF detected when $20^{\circ}<\Psi<30^{\circ}$). In the $Y_{\mathrm{GSM}}$ coordinate, the MMS apogee begins the magnetotail season at $Y_{\mathrm{GSM}}<0$, shifting to $Y_{\mathrm{GSM}}>0$ by the end of the season. Similarly, the apogee in $Z_{\mathrm{GSM}}$ starts positive and transitions to negative as the season progresses, in line with the tilt angle.
@@ 2.2 Solar wind and ground magnetic conditions
For determining the footpoints, it is necessary to incorporate data on solar wind and geomagnetic conditions. We used 1-minute resolution solar wind parameters, including solar wind dynamic pressure ($P_{dyn}$), interplanetary magnetic field (IMF), and solar wind velocity in GSE coordinates ($V_{i}$), sourced from OMNIweb https://omniweb.gsfc.nasa.gov/. The geomagnetic Dst index was replaced by the higher resolution Sym-h index, and the Kp index was replaced by the higher resolution Hp30 index (Yamazaki et al., 2022). Additionally, the solar wind driving parameters $G$ and $W$ (Qin et al., 2007) were obtained from the SpacePy Python package https://spacepy.github.io/index.html.
@@ 2.3 Mapping models
The footpoints for each of the 2079 BBF events were computed using the Tsyganenko models T89, T96, T01, and T04, available through the Geopack Python library (https://github.com/tsssss/geopack). Each model requires the event date and solar wind velocity to calculate the dipole tilt angle. The input parameters and valid regions for each model are summarized in Table 1.
Table 1: Kp and Dst: geomagnetic indices. $P_{dyn}$: the dynamic pressure of the solar wind; $B_{\mathrm{y,IMF}}$ and $B_{\mathrm{z,IMF}}$: y and z components of the IMF. G1 and G2: parameters related to solar wind and IMF conditions from the preceding 1-hour interval; W1 to W6: solar wind driving variables, quantifying the magnitudes of principal magnetospheric current systems. $r_{\mathrm{GEO}}$: geocentric distance
@@
T96 $-100<$ Dst $<20$ nT, $0.5\leq P_{\mathrm{dyn}}\leq 10$ nPa, $r_{\rm{GEO}}\leq 60R_{E}$
$|B_{\mathrm{y,IMF}}|\leq 10$ nT, $|B_{\mathrm{z,IMF}}|\leq 10$ nT
@@
In the case of T01 and T04 these models are only valid sunward of $X_{\rm{GSM}}=-15~{}R_{E}$. Consequently, only 12.5% of the 2079 BBF events satisfy this criterion and can be mapped to the Earth’s surface. Additionally, after visually inspecting the traced field lines, two criteria were applied to exclude events that could not be accurately modeled (see examples in Appendix A). The behavior of the models for such excluded events are beyond the scope of this work. The two criteria that were used to retain valid events are as follows:
@@
    The footpoint in the northern and southern hemisphere reaches the Earth’s surface.
@@
The database of 2079 BBF events with the Tsyganenko input and the footpoint output for each T89, T96, T01 and T04 model are available at Lanabere et al. (2024).
@@ 3 BBF footpoints from Different Tsyganenko Models
We analyzed the footpoint of the BBF events in both the Northern hemisphere and Southern hemisphere (NH and SH respectively) for each Tsyganenko model. The footpoint is presented in magnetic latitude (MLAT) and magnetic local time (MLT) in the altitude adjusted corrected geomagnetic (AACGM) coordinates. The altitude of the BBF footpoint is set at $462~{}\mathrm{km}$, which corresponds to the altitude of the Swarm satellites A and C.
@@ 3.1 Statistical Footpoint Positions for Each Tsyganenko Model
The footpoint histogram for each Tsyganenko model is shown in Figure 2. Statistically, the footpoints appear rather similar across all the mapping models, suggesting that despite the inherent differences in how these models represent Earth’s magnetosphere, they statistically converge on a similar spatial distribution of BBFs.
@@
Figure 2: footpoint histogram in MLAT and MLT (AACGM coordinates) of BBF in north and south hemisphere for different Tsyganenko models. Bin size: 1° x 0.5 hour. Altitude of footpoint is 462 km, about the same as Swarm A/C altitude.
@@
The magnetic latitude (MLAT) footpoints of BBFs are generally concentrated around $70^{\circ}\pm 4^{\circ}$ in both hemispheres. The spread in MLAT is related to the observed MMS location and could be indicative also of variations in geomagnetic conditions or differences in the way each Tsyganenko model maps to the ionosphere. In terms of magnetic local time (MLT), BBFs predominantly have their ionospheric footpoints around midnight ±$\pm 3$ hours, with a clear peak in the pre-midnight sector presumably associated with substorms.
@@
While the statistical distribution of BBF footpoints is similar across Tsyganenko models, significant variations can occur for specific events. These differences are crucial when searching for conjunctions. To better understand these differences, we perform a comparative analysis of the footpoint locations produced by each Tsyganenko model, using the T89 model as the reference. T89 was chosen as the benchmark because it maps the largest number of BBFs, thanks to its broader inclusion criteria and fewer restrictions on input parameters, although it is the least accurate due to its simpler design and reliance on a smaller dataset.
3.2 Comparison of BBF Footpoints: T89 vs. Later Models
@@
Figure 3(a-f) presents the MLT footpoint positions for each hemisphere, comparing the T89 model with T96, T01, and T04 models. Additionally, the mean bias error (MBE) and root mean square error (RMSE) are included to quantify the differences between the models.
@@
The MLT position between the models is quite similar, with differences typically within ±$\pm 1$ hour MLT. At $70^{\circ}$ latitude, this corresponds to a distance of approximately 571 km. However, from Figure 3a,d we notice that in the northern hemisphere, T96 tends to shift footpoints towards later MLT in the post-midnight sector and earlier MLT in the pre-midnight sector compared to the T89 footpoints. In contrast, the southern hemisphere shows the opposite behavior. The largest MLT differences are observed between T89 and T04 in the southern hemisphere.
@@
The MLAT footpoint position is shown in Figure 3(g-l) together with the MBE and RMSE. Compared to the MLT coordinate, MLAT differences between models are less pronounced in terms of distance. To put this into perspective, approximately $5^{\circ}$ in latitude corresponds to the same distance as 1 hour of MLT at $70^{\circ}$, as indicated by the pink dashed line. The MBE between T89 and T96 is $\mathrm{MBE}=0.22^{\circ}$ and $\mathrm{MBE}=0.81^{\circ}$ for northern and southern hemispheres respectively. This suggests that MLAT positions calculated using T96 are, on average, closer to the pole than those from T89. A similar trend, but with larger MBE values ($\mathrm{MBE}=1.18^{\circ}$ and $\mathrm{MBE}=1.14^{\circ}$ for northern and southern hemispheres), is observed between T89 and T01.
@@
Figure 3: Comparison of footpoint position in AACGM coordinates for T89 versus the T96, T01, and T04. (a-f) Magnetic local time (MLT). (g-l) Magnetic Latitude (MLAT). The black diagonal line represents the ideal scenario where both models agree perfectly. The red dashed lines indicate a deviation of ±$\pm 1$ (hour for the MLT plots and degree for MLAT plots). The pink dashed line marks a $5^{\circ}$ difference in latitude, corresponding to the approximate distance of 1 hour of MLT at $70^{\circ}$ latitude. The solid red line represents the mean value. For each plot the mean bias error (MBE) and the root mean square error (RMSE) are indicated.

### v3

@@ Abstract
Magnetospheric-ionospheric coupling studies often rely on multi-spacecraft conjunctions, which require accurate magnetic field mapping tools. For example, linking measurements from the magnetotail with those in the ionosphere involves determining when the orbital magnetic footpoint of THEMIS or MMS intersects with the footpoint of Swarm. The Tsyganenko models are commonly used for tracing magnetic field lines. In this study, we aim to analyze how the footpoint locations are impacted by the input parameters of these models, including solar wind conditions, geomagnetic activity, and the location in the magnetotail. A dataset of 2394 bursty bulk flows (BBFs) detected by MMS was mapped to Earth’s ionosphere with six different Tsyganenko models. Approximately 90% of the ionospheric footpoints are concentrated within $70^{\circ}\pm 5^{\circ}$ magnetic latitude (MLAT) and ±$\pm 3$ hours of magnetic local time (MLT) around midnight, with a pronounced peak in the pre-midnight sector. The MLT position showed a difference of approximately ±$\pm 1$ hour MLT across the models. Footpoint locations were linked to the dawn-dusk position of the BBFs, with differences between models associated with variations in the interplanetary magnetic field clock angle. The MLAT values exhibited similar differences of approximately ±$\pm 4^{\circ}$ around the mean value, with a systematic shift toward lower latitudes in the T89 model. This position is also influenced by the input parameters of the model representing the dynamics of Earth’s magnetosphere, where stronger magnetospheric activity typically corresponds to lower latitudes. The uncertainty on the BBF footpoint location impacts the number of conjunctions with Swarm. Generally, Swarm B exhibited more conjunctions than Swarm A or C in the Northern Hemisphere. However, when considering only Swarm-BBF conjunctions where the distance between footpoints computed with T89 and TA15n is smaller than the size of the BBF footprint, the number of conjunctions is reduced to less than half of the total.
@@
Bursty bulk flows footpoint – Tsyganenko models – Field line tracing – BBF footpoint-Swarm Conjunctions
@@ 1 Introduction
Although model accuracy (i.e., how closely the derived footpoint matches its true value) is not the focus of this study, previous research has highlighted limitations in these models. For instance, Pulkkinen and Tsyganenko (1996) analyzed the accuracy of the T89 model, finding that the mapping errors were largest on the dayside, while on the nightside they ranged from $0.5\degree$ to $1.8\degree$ in latitude (63°72°63°72°63$\degree$–72$\degree$63 ° – 72 °), increasing with geomagnetic activity. Similarly, Shevchenko et al. (2010) reported errors of $\sim 1\degree$ when evaluating the accuracy of the T96 model. Thus, the T96 model performs a better mapping than T89, but in both cases the mapping error increases with magnetic activity. These discrepancies have been observed in case studies. For instance, Opgenoorth et al. (1994) found that a localized FAC was observed a few degrees north compared to the location obtained with the T89 model. Similarly, Wei et al. (2021) noted differences in the ground track mapping when using three different Tsyganenko models (T87, T89, and T96) while investigating the characteristics and responses of the magnetosphere-ionosphere-ground system during a case study.
This study utilizes a large database of BBF events observed by the Magnetospheric Multiscale (MMS) spacecraft (Burch et al., 2016) to statistically compare the mapped footpoint locations among six different Tsyganenko models. We refer to the footpoint as a unique point representing the BBF and do not consider the spatial size of the BBF mapped into the ionosphere (footprint). In this work, we do not address the error in the location of the footpoint, since this implies establishing a ground truth for the correct footpoint for each event, which is beyond the scope of this analysis. Instead, we analyze how the footpoint locations are impacted by the input parameters of these models, including solar wind conditions, geomagnetic activity, the location of the BBF in the magnetotail, and the tilt angle. We report how the number of conjunctions between BBFs and Swarm spacecraft changes when the conjunction definition is adjusted. Specifically, we examine how increasing the time interval between the detection of a BBF and the Swarm satellites’ path through/nearby the BBF footpoint, as well as varying the footpoint magnetic latitude error, affects the occurrence and quality of conjunctions. Although we did not conduct a study on the error in the location, we hope that the quality indicator will help researchers determine when to be more cautious when studying conjunctions.
@@ 2.1 The BBF database
Since our goal is to find the ionospheric footpoint of the BBFs, we restricted our database to Earthward-moving BBFs, which constitute 89% of the total database. The mean and median duration of these Earthward BBFs are 3.60 minutes and 2.63 minutes, respectively, with approximately 80% of the events lasting less than 5 minutes. After excluding an outlier with a duration of $\approx 3~{}$hours, the BBF dataset consists of 2135 events, with the longest event lasting 35 minutes.
@@
Additionally, Figure 1b-c show the MMS orbit for the second MMS season (2018-05-25 to 2018-09-28) and the BBFs detected during the same period (blue dots correspond to BBF detected when $20^{\circ}<\Psi<30^{\circ}$). In the $Y_{\mathrm{GSM}}$ coordinate, the MMS apogee begins the magnetotail season at $Y_{\mathrm{GSM}}<0$, shifting to $Y_{\mathrm{GSM}}>0$ by the end of the season. Similarly, the apogee in $Z_{\mathrm{GSM}}$ starts positive and transitions to negative as the season progresses, in line with the tilt angle.
@@ 2.2 Solar wind and ground magnetic conditions
For determining the footpoints, it is necessary to incorporate data on solar wind and geomagnetic conditions. We used 1-minute resolution solar wind parameters, including solar wind dynamic pressure ($P_{dyn}$), interplanetary magnetic field (IMF), and solar wind velocity in GSE coordinates ($V_{i}$), as extracted from NASA/GSFC’s OMNI data set through OMNIWeb. The geomagnetic Dst index was replaced by the higher resolution Sym-h index, and the Kp index was replaced by the higher resolution Hp30 index (Yamazaki et al., 2022). Additionally, the solar wind driving parameters $G$ and $W$ (Qin et al., 2007) have been used as input parameters for T01 and T04 respectively, as well as the solar wind coupling functions B-index (Boynton et al., 2011) and N-index (Newell et al., 2007) for the TA15b and TA15n models, respectively.
@@ 2.3 Mapping models
The footpoints for each of the 2079 BBF events were computed using the Geopack tracing routine for the Tsyganenko models T89, T96, T01, and T04, available through the Geopack Python library [^1] https://github.com/tsssss/geopack, and TA15b, TA15n, available through the IDL Geopack DLM [^2] https://korthhaus.com/idl-software/idl-geopack-dlm/. Each model requires the event date and solar wind velocity to calculate the dipole tilt angle. The input parameters and valid regions for each model are summarized in Table 1.
Table 1: Kp and Dst are geomagnetic indices; $P_{dyn}$ the dynamic pressure of the solar wind; $B_{\mathrm{y,IMF}}$ and $B_{\mathrm{z,IMF}}$ the y and z components of the IMF; G1 and G2 are the parameters related to solar wind and IMF conditions from the preceding 1-hour interval; W1 to W6 are the solar wind driving variables, quantifying the magnitudes of principal magnetospheric current systems; B-Index and N-index are solar wind coupling function, suggested by Boynton et al. (2011) and Newell et al. (2007) respectively. $r_{\mathrm{GEO}}$ is the geocentric distance
@@
T96 Dst,$P_{\mathrm{dyn}}$ nPa, $B_{\mathrm{y,IMF}}$, $B_{\mathrm{z,IMF}}$, $r_{\rm{GEO}}\leq 60R_{E}$
@@
TA15b $P_{dyn}$, $B_{\mathrm{y,IMF}}$, $B_{\mathrm{z,IMF}}$, $B\mathrm{-Index}$ $r_{\rm{GEO}}\leq 60R_{E}$
TA15n $P_{dyn}$, $B_{\mathrm{y,IMF}}$, $B_{\mathrm{z,IMF}}$, $N\mathrm{-Index}$ $r_{\rm{GEO}}\leq 60R_{E}$
The valid ranges for each input parameter are provided in the respective documentation for each model. However, we emphasize that the T01 and T04 models are only valid sunward of $X_{\rm{GSM}}=-15~{}R_{E}$. Consequently, only 12.5% of the 2079 BBF events satisfy this criterion and can be mapped to the Earth’s surface. Additionally, after visually inspecting the traced field lines, two criteria were applied to exclude events that could not be accurately modeled (see examples in the Appendix A). The behavior of the models for such excluded events is beyond the scope of this work. The two criteria that were used to retain valid events are as follows:
@@
    The footpoint in the Northern and Southern Hemispheres reaches the Earth’s surface.
@@
The database of 2079 BBF events, including the Tsyganenko model inputs and the resulting magnetic footpoints in both the Northern and Southern Hemispheres for the T89, T96, T01, T04, TA15b, and TA15n models, is available in Lanabere et al. (2025). The footpoints at the Earth’s ionosphere are given in both GSM coordinates and the altitude-adjusted corrected geomagnetic (AACGM) coordinate system.
@@ 3.1 Statistical footpoint positions for each Tsyganenko model
We analyzed the footpoint of the BBF events in both the Northern Hemisphere and Southern Hemisphere (NH and SH respectively) for each Tsyganenko model. The footpoint histogram for each Tsyganenko model is shown in Figure 2. Statistically, the footpoints appear rather similar across all the mapping models, suggesting that despite the inherent differences in how these models represent Earth’s magnetosphere, they statistically converge on a similar spatial distribution of BBFs.
@@
Figure 2: BBF ionospheric footpoint histograms in MLAT and MLT (AACGM coordinates) for different Tsyganenko models in the Northern and Southern Hemispheres. Bin size: $1^{\circ}$ x 0.5 hour.
@@
The magnetic latitude (MLAT) footpoints of BBFs are generally concentrated around $70^{\circ}\pm 5^{\circ}$ in both hemispheres. The wider MLAT spread observed in the TA15 model may be attributed to its capability to represent the magnetosphere under a wider variety of solar wind conditions. In terms of magnetic local time (MLT), BBFs predominantly have their ionospheric footpoints around midnight ±$\pm 3$ hours, with a clear peak in the pre-midnight sector presumably associated with substorms.
@@
While the statistical distribution of BBF footpoints is similar across the Tsyganenko models, significant variations can occur for specific events. These differences are crucial when searching for conjunctions. To better understand these differences, we performed a comparative analysis of the footpoint locations produced by each Tsyganenko model, using the TA15n model as the reference. TA15n was chosen as the benchmark because it maps a large number of BBFs, due to its broader inclusion criteria, and is expected to be more accurate due to its more sophisticated representation and larger dataset.
3.2 Comparison of BBF footpoints: TA15 vs. previous models
@@
The MLT footpoint positions for each hemisphere are shown in Figure 3(a–j), comparing the TA15n model with the T89, T96, T01, T04, and TA15b models. Additionally, the mean bias error (MBE) and the root mean square error (RMSE) are included to quantify the differences between the models.
@@
The MLT positions between the models are quite similar, with differences typically within ±$\pm 1$ hour MLT. At $70^{\circ}$ latitude, this corresponds to a distance of approximately 571 km. However, as shown in Figure 3a,f T89 tends to shift footpoints toward earlier MLT in the post-midnight sector and toward later MLT in the pre-midnight sector, relative to the TA15n footpoints in both hemispheres. A similar, though less pronounced, trend is observed between T96 and TA15n in the Southern Hemisphere, while the opposite trend is observed in the Northern Hemisphere.
@@
The MLAT footpoint position is shown in Figure 3(k-t) together with the MBE and RMSE. The differences between models in MLAT are of the same order of magnitude as those in MLT. To put this into perspective, approximately $5^{\circ}$ in latitude corresponds to the same distance as 1 hour of MLT at $70^{\circ}$, as indicated by the pink dashed line. In general, the TA15 models show a wider spread in MLAT than older models. The MBE between TA15n and T89 is $\mathrm{MBE}=1.03^{\circ}$ and $\mathrm{MBE}=1.21^{\circ}$ for the Northern and Southern Hemispheres, respectively, suggests that the MLAT positions calculated using TA15n are, on average, $1^{\circ}$ closer to the pole than those from T89. A similar poleward bias is observed when comparing TA15n to T96, though with smaller MBE values ($\mathrm{MBE}=0.57^{\circ}$ and $\mathrm{MBE}=0.44^{\circ}$ for the Northern and Southern Hemispheres). In contrast, the comparison between TA15n and T01 shows the opposite trend, with T01 footpoints located closer to the pole on average. No consistent latitudinal bias is observed between TA15n and T04. Due to the broader spread in MLAT for the TA15 model compared to older models, the RMSE values show that the MLAT footpoint location derived from TA15 and from (T89,T96, T01, T04) disagree by about $2^{\circ}$ of latitude on average, with the largest discrepancies observed when compared to T89.
@@
Figure 3: Comparison of footpoint position in AACGM coordinates for Ta15n versus the T89, T96, T01, T04, and TA15b. (a-j) Magnetic local time (MLT). (k-t) Magnetic Latitude (MLAT). The black diagonal line represents the ideal scenario where both models agree perfectly. The red dashed lines indicate a deviation of ±$\pm 1$ (hour for the MLT plots and degree for MLAT plots). The pink dashed line marks a $5^{\circ}$ difference in latitude, corresponding to the approximate distance of 1 hour of MLT at $70^{\circ}$ latitude. The solid red line represents the mean value. For each plot the mean bias error (MBE) and the root mean square error (RMSE) are indicated.